OpenStudy (thediamondheart):

Simplify. –1 + 5(6 – y) A. –y + 5 B. –y + 29 C. –5y + 29 D. –5y – 31

OpenStudy (bhumi273):

-1+5(6-y)=0 5(6-y)=1 5*6-5*y=1 30-5y=1 30-1-5y=0 29-5y=0 -5y+29 therefore ans. is (c)